# Robustness-aware Automatic Prompt Optimization

### v2

@@ 4.3 Adversarial Attack Phase
Specifically, for the two distinct perturbation types, P1 and P2, we draw on the idea of Tramer and Boneh (2019) and design two different modes of adversarial attacks. This ensures that the text used in the adversarial training process encompasses all perturbations within each type. As a result, the generated prompts demonstrate robustness against all attacks corresponding to their respective types. For the P1 type, we propose a mix-mode, which is implemented as follows:
@@
Where $\mathcal{P(\cdot)}$ denotes that the given input $x$ is perturbed under the guide $g$, $g_{n}$in P1, $x_{n}$ denotes the text generated under each perturbation. The rationale behind this design is that, under the perturbation of the P1 mode, the superposition of each perturbation has minimal impact on the effect of other perturbations. For example, the input "I like apple" becomes "I lide apple" after the C1 perturbation, and then the output becomes "I lide apple.@jjs" after the C3 perturbation. This design enables adversarial attacks to generate results that effectively incorporate all types of perturbations. Thus, the robustness of the generated results against all types of perturbations in P1 can be ensured, while simplifying the input for the second stage.
@@
where $\mathcal{U(\cdot)}$ denote the set generated after all perturbations. Unlike P1 type perturbations, which do not interfere with each other’s output, different perturbations in P2 type can affect the results of other perturbations. Therefore, we choose to combine them and input them into the adversarial optimization phase to train the robustness of the generated prompts against all types of perturbations.
4.4 Adversarial Optimization Phase
In the optimization stage, we utilize the adversarial sample $x^{\prime}$ generated in the first stage. We then analyze the differences between $x^{\prime}$ and the corresponding original text $x$, and use these differences to compute the gradient $g^{\prime}$ when the guide is used as the optimization prompt. The detailed process is as follows:
@@ 4.4 Adversarial Optimization Phase
Where $\mathcal{D}(\cdot)$ denotes the generating difference and $\mathcal{G}(\cdot)$ denotes the generation of general guidence. In the iterative optimization process following gradient generation, we also use the prompt’s score on the task as the loss function. However, unlike the previous stage, we select the prompt with the highest score at each iteration to maximize task performance. To ensure that the final prompt is robust to all perturbations, we calculate the score for each perturbation across the vulnerabilities of the target task. The specific loss function is as follows:
@@
Methods C2 perturbation C1 perturbation S1 perturbation
Rouge-1 Rouge-2 Rouge-L Rouge-1 Rouge-2 Rouge-L Rouge-1 Rouge-2 Rouge-L
DA 11.37 2.65 9.78 11.90 2.48 9.93 10.90 2.38 9.11
MI 14.55 2.80 12.31 15.27 3.12 12.80 15.13 3.19 12.65
EvoPrompt 17.62 3.16 14.96 17.24 3.03 14.61 17.71 3.24 15.34
BATprompt 18.31 3.08 15.50 18.65 2.91 15.03 17.65 2.81 15.58
BATprompt 21.10 4.36 16.85 21.17 4.37 16.34 21.47 4.59 16.78
Table 2: The Rouge-1($\uparrow$) score, Rouge-2($\uparrow$) score and Rouge-L($\uparrow$) score obtained by the prompts generated by the three methods for the task where the text summarization task is weak under the P1 class perturbation.
DA 12.18 2.87 10.25 11.85 2.92 10.46 12.62 2.66 10.60 12.12 2.91 10.25
MI 15.55 3.15 13.01 15.09 3.19 12.88 14.06 3.03 11.69 15.29 2.74 13.13
EvoPrompt 17.54 3.09 14.93 17.20 3.03 14.85 16.97 2.67 14.48 17.03 2.93 14.49
BATprompt 18.42 3.10 15.82 18.64 3.39 15.98 17.92 2.91 15.28 17.90 3.06 15.46
BATprompt 22.04 4.91 17.29 22.09 4.71 16.70 21.51 4.76 16.45 21.97 4.55 16.78
Table 3: The Rouge-1($\uparrow$) score, Rouge-2($\uparrow$) score and Rouge-L($\uparrow$) score obtained by the prompts generated by the three methods for the task where the text summarization task is weak under the P2 class perturbation.
@@ 5.1 Implementation Details and Baselines
In the experiments, we use GPT-3.5-turbo to do the adversarial training and use GPT-3.5-turbo, GPT-4o-mini and Llama2-7b to test the effectiveness of the instructions generated BATprompt. For perturbations of type P1, we select five examples in each iteration, while for perturbations of type P2, we select three examples per iteration, considering the number of adversarial examples generated. After five iterations, we choose the prompt with the highest score on the training set and evaluate its performance on the test set. In the evaluation, we compare the prompts generated by BATprompt with the following methods:
  • EvoPrompt Guo et al. (2023): Evoprompt uses a genetic algorithm to optimize prompts, its tasks contain both language understanding and language generation tasks.
@@
  • Manual Instructions(MI): These instructions are based on existing work that is task-specific guidelines. Including language understanding task Zhang et al. (2023a), text simplification task Zhang et al. (2023b) and summarization task Sanh et al. (2021).
  • Natural Instructions(NI) Mishra et al. (2021): This contains manually designed prompts across a diverse range of datasets and tasks.
@@
  • Data Augmentation(DA): We use the data augmentation method as baseline, which takes the perturbed text as input data and iteratively optimizes it, to explore whether this method can remain robust to all types of text perturbations.
5.2 Data Generation and Metrics
To evaluate the effectiveness of BATprompt, we construct datasets with various perturbations tailored to different tasks. Using an iterative approach guided by gradients, we introduce perturbations into the datasets. However, different from the traditional method of adversarial attacks, which searches attacks freely in all ranges, our method adds attacks in specific perturbation spaces. For dataset selection, in the language understanding tasks, we focus on six datasets to apply and test perturbations. These include sentiment classification datasets: SST-2 Socher et al. (2013), CR Hu and Liu (2004), SST-5 Socher et al. (2013), and MR PaNgB (2005), as well as topic classification datasets: AG’s News Zhang et al. (2015) and TREC Voorhees and Tice (2000) with the Prediction accuracy as the score. In language generation tasks, we utilize the Asset Alva-Manchego et al. (2020) for text simplification, which includes multiple benchmarks for reference translations. In this task we use SARI Xu et al. (2016) as the metrics which is an n-gram-based scoring system extensively utilized for text editing tasks. For the text summarization task, we use the XSum Narayan et al. (2018), which consists of concise summaries generated from longer texts. In this task, we use Rouge-1, Rouge-2, Rouge-L as metrics Lin (2004), which widely used to evaluate the quality of generated text tasks. They focus on the number of n-grams with the same outcome and the overlap of the longest common subsequence.
5.3 Effectiveness on Language Generation
In this section, we utilize the GPT-3.5-turbo model to evaluate the generated prompts. For the text summarization task, we assessed the quality of the generated prompt guidance using three metrics, ROUGE-1, ROUGE-2, and ROUGE-L, calculated on the test set. For text simplification tasks, the quality of the generated prompts were by calculating the SARI score of the prompt guided result.
For the text summarization task, the results for the perturbations P1 and P2 are presented in  Table 2 and  Table 3. Compared to previous prompts and their generation methods, the prompts generated by BATprompt demonstrate a significant performance improvement on the perturbed datasets. Furthermore, BATprompt exhibits strong robustness across all types of perturbations, achieving superior Rouge-1, Rouge-2, and Rouge-L scores compared to other prompts in all aspects. Notably, BATprompt outperforms the second-best method by an impressive 23% under C2 perturbations.
For the text simplification task, the SARI values achieved by BATprompt under P1 and P2 perturbations are shown in the  Table 4. Similar to the text summarization task, BATprompt demonstrates strong robustness across all perturbations. Its SARI values consistently have better performance than those of the second-best method, highlighting its effectiveness and resilience.
DA 44.87 42.53 44.04 44.89 44.66 45.28
MI 36.67 34.77 37.43 37.18 35.93 37.23
EvoPrompt 44.61 44.89 45.39 47.38 46.07 49.09
BATprompt 45.28 43.45 44.53 46.29 46.47 47.90
BATprompt 45.79 45.11 49.50 47.74 47.35 49.55
Table 4: The SARI($\uparrow$) score obtained by the prompts generated by the two methods for the task where the text simplification task.
Type Methods CR SST-2 MR SST-5 TREC AG’news Avg.
P1 DA 88.5 89.8 85.5 44.0 43.8 86.3 73.0
MI 89.3 88.8 85.5 45.5 50.3 87.0 74.4
NI 83.3 80.3 80.0 23.3 45.3 65.3 62.9
EvoPrompt 63.8 80.5 78.1 10.5 1.5 67.3 51.0
BATprompt 85.5 86.0 81.0 45.3 49.8 85.5 72.1
BATprompt 89.8 90.0 86.3 46.0 52.3 88.5 75.4
P2 DA 85.1 85.5 81.2 40.1 50.3 86.7 71.5
MI 85.0 85.5 82.0 43.6 46.2 86.2 71.4
NI 79.1 80.0 77.0 20.0 43.4 61.8 60.2
EvoPrompt 62.8 78.1 80.6 8.6 1.5 66.8 49.7
BATprompt 80.6 81.5 78.1 40.9 49.9 84.3 69.2
BATprompt 85.7 86.3 82.8 44.9 51.6 87.7 73.2
Table 5: Average score($\uparrow$) of the prompts from different method on six language understanding datasets. The table in the upper half is a perturbation of type P1, and the table in the lower half is a perturbation of type P2.
Figure 4: The performance of the prompts generated by BATprompt on undisturbed datasets of language understanding task.
5.4 Effectiveness on Language Understanding
In this part, we use GPT-3.5-turbo to evaluate the accuracy of its judgments on six datasets, which is recorded as its score. For each dataset, we train the model on perturbations from class P1 and perturbations from class P2. We then test the optimized prompts across different types of perturbations and compute the average performance over them. The results for the perturbations from class P1 and class P2 are presented in the  Table 5. Specific experimental indicators are given in Figure 12.
The results demonstrate that the prompts generated by BATprompt outperform existing prompts across the six text understanding datasets, exhibiting notable stability against the seven perturbations to which BATprompt itself is not inherently robust. Its performance is particularly impressive. Notably, BATprompt achieves a 3% improvement on the TREC dataset for P1 perturbations and a remarkable 12% improvement for T2 perturbations.
5.5 Model Transferbility and Perturbation Transferbility
We also use GPT-4o-mini and Llama2-7b to test the effectiveness of our method. The results are showed in the subsection D.3. To further assess the robustness of our prompts against unseen perturbations, we evaluated them on perturbations that were not included in the training set. As shown in subsection D.4, although their performance is marginally lower than that of prompts specifically optimized for those perturbations, our approach consistently outperforms all baseline methods.
@@ 5.6 Data Augmentation
To evaluate the limitations of the data augmentation method, we exclude the use of BATprompt during testing. Instead, we treat all the perturbed data as the training set and allow the LLM to generate prompt based on this training set. The results for three different tasks are presented in the DA column across the  Table 4Table 2,  Table 3 and  Table 5. As shown, the prompts generated using the data augmentation method lack robustness to perturbations across all three tasks. Furthermore, their performance is even worse than unoptimized prompts when tested on handwritten perturbations. We attribute this degradation to the excessive diversity of perturbations, which hinders the LLM’s ability to focus accurately on the perturbations, leading to a decline in performance.
5.7 Albation Study
@@
To evaluate the effectiveness of adversarial training strategies for prompt optimization, we removed the adversarial training part from BATprompt and used only a simple iterative optimization method. This allowed us to isolate and assess the impact of adversarial training. In the  Table 2,  Table 3,  Table 4 and  Table 5, BATprompt represents the specific scores of the prompts generated using only the iterative optimization method for both the text comprehension and text generation tasks. The experimental results demonstrate that, across all tasks, prompts generated purely through iterative optimization perform worse than those generated by BATprompt across all types of perturbations. This proves the effectiveness of our adversarial training strategy. In addition, we also explore the influence of the number of training iterations on the quality of the generated prompt, and the experimental results show that the results are slightly higher when the number of iterations is 5 than other iterations. The specific result is in subsection E.1.

### v3

@@ 3.3 Pseudo Gradient Perturbation Phase
At this stage, we guide the construction of the perturbation dataset through the perturbation category, and maintain the similarity as the purpose of building the benchmark. However, for the two distinct perturbation types, P1 and P2, we design two different modes of perturbation. This ensures that the text used in the optimization process encompasses all perturbations within each type. As a result, the generated prompts demonstrate robustness against all attacks corresponding to their respective types. For the P1 type, we propose a mix-mode, which is implemented as follows:
@@
Where $\mathcal{P(\cdot)}$ denotes that the given input $x$ is perturbed under the guide $g$, $g_{n}$ $\in$ P1, $x_{n}$ denotes the text generated under each perturbation. After each perturbation, we compute the Levenshtein distance between all perturbed texts and their corresponding source texts, selecting the resulting text as input for the next iteration. The rationale behind this design is that, under the perturbation of the P1 mode, the superposition of each perturbation has minimal impact on the effect of other perturbations. This design enables to generate results that effectively incorporate all types of perturbations.
@@
where $\mathcal{U(\cdot)}$ denote the set generated after all perturbations. Unlike P1 type perturbations, P2 perturbations involve semantic modifications through operations such as rewriting, changing grammar and so on. To ensure that the meaning remains aligned with the original sentence, thus preventing any adverse impact on the performance of the LLMs. We rely on semantic similarity as the evaluation criterion. We choose to combine and input them into the pseudo gradient optimization phase.
Methods P1 perturbation P2 perturbation P1 perturbation P2 perturbation
Rouge-1 Rouge-2 Rouge-L Rouge-1 Rouge-2 Rouge-L SARI SARI
DA 11.39 2.50 9.61 12.19 2.84 10.39 44.87 44.28
MI 14.98 3.04 12.59 15.00 3.03 12.68 36.67 36.51
EvoPrompt 17.52 3.14 14.97 17.19 2.93 14.69 44.61 46.56
PGO 18.20 2.93 15.37 18.22 3.12 15.64 45.28 45.73
PGO 21.25 4.44 16.66 21.90 4.73 16.81 45.79 47.85
Table 3: The average Rouge-1($\uparrow$) score, Rouge-2($\uparrow$) score and Rouge-L($\uparrow$) score on the text summarization task and average SARI($\uparrow$) scores on the text simplification task for our method and the comparison methods.
3.4 Pseudo Gradient Optimization Phase
In the optimization stage, we utilize the perturbed sample $x^{\prime}$ generated in the first stage. We then analyze the differences between $x^{\prime}$ and the corresponding original text $x$, and use these differences to compute the gradient $g^{\prime}$ when the guide is used as the optimization prompt. The detailed process is as follows:
@@ 3.4 Pseudo Gradient Optimization Phase
Where $\mathcal{D}(\cdot)$ denotes the generating difference and $\mathcal{G}(\cdot)$ denotes the generation of general guidance. In the iterative optimization process following gradient generation, we also use the prompt’s score on the task as the loss function. However, unlike the previous stage, we select the prompt with the highest score at each iteration to maximize task performance. To ensure that the final prompt is robust to all perturbations, we calculate the score for each perturbation across the vulnerabilities of the target task. The specific loss function is as follows:
@@ 4.1 Implementation Details and Baselines
In the experiments, we use GPT-3.5-turbo to do the optimization training and use GPT-3.5-turbo, GPT-4o-mini and Llama2-7b to test the effectiveness of the instructions generated PGO. For perturbations of type P1, we select five examples in each iteration, while for perturbations of type P2, we select three examples per iteration, considering the number of perturbed examples generated. After five iterations, we choose the prompt with the highest score on the training set and evaluate its performance on the test set. More experiment settings in the Section  B.1. In the evaluation, we compare the prompts generated by PGO with the following methods:
@@
  • Manual Instructions(MI) and Natural Instructions(NI) Mishra et al. (2021): MI are based on existing work that is task-specific guidelines. Including classification task Zhang et al. (2023a), text simplification task Zhang et al. (2023b) and summarization task Sanh et al. (2021). NI contains manually designed prompts across a diverse range of datasets and tasks.
  • APE Zhou et al. (2022), InstructZero Chen et al. (2023), INSTINCT Lin (2004): These papers leverages the reasoning ability of LLMs to generate prompts based on both the inputs and the corresponding answers.
  • EvoPrompt Guo et al. (2023): Evoprompt uses genetic algorithms to optimize prompts for classification and generation tasks.
@@
  • Data Augmentation(DA) and PGO: DA uses the data augmentation method as baseline, which takes the perturbed text as input data and iteratively optimizes it, to explore whether this method can remain robust to all types of text perturbations. PGO uses the simple iterative method to optimize the prompts.
We also analysis the performance of adding suggestive suffix without PGO directly in Section  C.4 and the cost of PGO in Section  C.5. Our optimal results are shown in Appendix F.
4.1.1 Metrics
In the classification tasks, we use the prediction accuracy as the score. In language generation tasks, we use SARI Xu et al. (2016) as the metrics which is an n-gram-based scoring system extensively utilized for text editing tasks. For the text summarization task, we use Rouge-1, Rouge-2, Rouge-L as metrics Lin (2004), which widely used to evaluate the quality of generated text tasks. They focus on the number of n-grams with the same outcome and the overlap of the longest common subsequence.
Methods P1 perturbation P2 perturbation Avg.
CR SST-2 MR SST-5 TREC AG CR SST-2 MR SST-5 TREC AG
DA 88.5 89.8 85.5 44.0 43.8 86.3 85.1 85.5 81.2 40.1 50.3 86.7 72.3
MI 89.3 88.8 85.5 45.5 50.3 87.0 85.0 85.5 82.0 43.6 46.2 86.2 72.9
NI 83.3 80.3 80.0 23.3 45.3 65.3 79.1 80.0 77.0 20.0 43.4 61.8 61.6
EvoPrompt 63.8 80.5 78.1 10.5 1.5? 67.3 62.8 78.1 80.6 8.6 1.5? 66.8 50.4
PGO 85.5 86.0 81.0 45.3 49.8 85.5 80.6 81.5 78.1 40.9 49.9 84.3 70.8
PGO 89.8 90.0 86.3 46.0 52.3 88.5 85.7 86.3 82.8 44.9 51.6 87.7 73.2
Table 4: Average score($\uparrow$) of the prompts from different method on six classification datasets. ?? Indicates that we provide specific analysis of the value of this indicator in Section  C.3. AG indicates AG’news dataset.
@@ 4.2 Data Augmentation
To evaluate the limitations of the data augmentation method, we exclude the use of PGO during testing. Instead, we treat all the perturbed data as the training set and allow the LLMs to generate prompt based on this training set. The results for three different tasks are presented in the DA row across the Table. 3, Table. 4. As shown, the prompts generated using the data augmentation method lack robustness to perturbations across all three tasks. Furthermore, their performance is even worse than unoptimized prompts when tested on handwritten perturbations. We attribute this degradation to the excessive diversity of perturbations, which hinders the LLMs’s ability to focus accurately on the perturbations, leading to a decline in performance.
4.3 Experiments Results
In this part, we use GPT-3.5-turbo to test the performance of prompts generated by PGO and other comparison methods on the three tasks. The results of text summarization task and text simplification task are shown in the Table 3. Compared to previous prompts and their generation methods, the prompts generated by PGO demonstrate a significant performance improvement on the perturbed datasets. It is worth noting that in the text summarization task, the model outperforms the second-best baseline by an average of 20%, 46%, and 7.5% in ROUGE-1, ROUGE-2, and ROUGE-L scores, respectively. For classification task, the results for the perturbations from class P1 and class P2 are presented in the Table 4. The table demonstrates that the prompts generated by PGO outperform existing prompts across the six text understanding datasets, exhibiting notable stability against the seven perturbations to which PGO itself is not inherently robust. It is worth mentioning that PGO achieves a 4% improvement on TREC under the P1-class perturbation and a 3% improvement on SST-5 under the P2-class perturbation. In addition, we observe that the cue words generated by our method demonstrate notably stronger retention robustness on multiple-choice datasets such as SST-5, TREC, and AG News. We note that EvoPrompt performs significantly worse than our method on the SST-5 and TREC datasets, and We attribute this to deficiencies in the generated prompts resulting from its prompt optimization method; a detailed analysis can be found in the Section  C.3. For methods APE Zhou et al. (2022) InstructZero Chen et al. (2023), INSTINCT Lin (2004), PGO still maintains the optimal performance on all tasks, and more specific results are in the Section  C.1.
4.4 Transferability and Generalization
4.4.1 Model Transferability
To evaluate the transferability of our generated prompts, we select the text summarization task and assess the effectiveness of our method on GPT-4o-mini and LLaMA2-7B. The results demonstrate that, compared to other methods, our approach consistently maintains strong performance across different models. Additional transferability experiments are provided in the Section  C.6.
4.4.2 Perturbation Generalization
In this section, we evaluate the transferability of prompts trained under different perturbation types. Specifically, we adopt a cross-testing strategy: prompts trained under perturbation P1 are tested on robustness against perturbation P2, and vice versa. The results are presented in the Table. 6. Notably, our PGOchange performs only slightly worse than PGO across most metrics, indicating that prompts optimized for one class of perturbations can still exhibit strong robustness when applied to other perturbation types.
Model Method Rouge-1 Rouge-2 Rouge-L
GPT-4o-mini MI 16.10 2.84 13.00
EvoPrompt 16.83 2.63 14.33
PGO 19.49 3.63 14.97
LLaMA2-7b MI 16.45 2.95 13.80
EvoPrompt 14.80 2.56 12.85
PGO 18.30 3.64 14.71
Table 5: Transferability of prompts across methods and models in text summarization.
4.5 Ablation Study
4.5.1 The Effectiveness of Pseudo Gradient
@@
To evaluate the effectiveness of first stage in the prompt optimization, we removed the pseudo gradient perturbation part from PGO and used only a simple iterative optimization method. This allowed us to isolate and assess the impact of first phase. The results are show in PGO rows in the Table 3 and Table. 4. The experimental results demonstrate that, across all tasks, prompts generated purely through iterative optimization perform worse than those generated by PGO across all types of perturbations. This proves the effectiveness of our pseudo gradient strategy.
4.5.2 Iteration Numbers
In our optimization process, we set the number of iterations to five, as we observed that the prompts optimized by the LLMs achieved optimal performance at the fifth iteration. To illustrate this, we used the text summarization task as an example. We tested the performance of prompts generated across six iterations (from 1 to 6), and the results are presented in the Figure. 5. For each iteration, we evaluate the prompts across different perturbation types and calculate their average performance. The results show a consistent improvement in prompt performance during the initial rounds, with the optimal performance observed around the fourth or fifth iteration. This trend suggests that the LLMs effectively refines the prompts through iterative gradient-guided optimization and semantic space exploration, progressively approaching the optimal solution. However, beyond the fifth iteration, the performance of the prompts have a slightly declines. In order to speed up the inference experiment and reduce the cost, we choose the number of iterations to be 5. We present more discussion in Appendix D.
Type Method Rouge-1 Rouge-2 Rouge-L
P1 PGOchange 21.35 4.59 16.10
PGO 21.25 4.44 16.66
P2 PGOchange 21.26 4.31 16.33
PGO 21.90 4.73 16.81
Table 6: The average of Rouge-1($\uparrow$), Rouge-2($\uparrow$) and Rouge-L($\uparrow$) of prompts trained by different perturbation on the text summarization task.PGOchange means test the results on other perturbation.
Figure 5: Relationship between iteration count and PGO prompt performance.
5 Related Work
5.1 Prompt Optimization
Manually crafted prompts often fail to improve LLM performance, motivating the rise of prompt optimization.  Li et al. (2023) introduces a fine-tuning approach based on context and probability ordering, while reinforcement learning methods such as Eureka Ma et al. (2023), Prompt-OIRL Sun et al. (2023), and RLPrompt Deng et al. (2022) further enhance optimization. Recent work leverages LLMs themselves for prompt generation and refinement, including APE Zhou et al. (2022), self-correction via pseudo gradients Pryzant et al. (2023), genetic algorithms Guo et al. (2023), and robustness studies Jin et al. (2024); Zhou et al. (2024a). However, most methods rely on clean data, neglecting natural input perturbations that degrade performance. Our approach addresses this by emphasizing robustness against such perturbed inputs.
5.2 Improving the Robustness of LLMs
Recent studies explore adversarial attacks on LLMs, where slight input perturbations (e.g., spelling errors) mislead models into incorrect outputs.  Zhou et al. (2024c) target math reasoning, while  Zhu et al. (2023) categorize hint attacks into four types, detailed by  Xu et al. (2024). Common approaches include word-level Wang et al. (2023a); Zhou et al. (2024b) and sentence-level Gu et al. (2023); Dong et al. (2021) attacks.  Miao et al. (2024) introduce attacks for T2M, and  Raina et al. (2024) reveal Judge-LLMs’ vulnerability, inflating evaluation scores. Other works Yao et al. (2023); Kumar et al. (2023) study adversarial behaviors and propose defense strategies, including adversarial hallucination Kumar et al. (2023) and refusal-based defenses Sheshadri et al. (2024); Xhonneux et al. (2024); Lin and Zhao (2024). Adversarial training methods like ReFAT Sheshadri et al. (2024), Latent Adversarial Training Xhonneux et al. (2024), and LLAMOS Lin and Zhao (2024) further enhance LLM robustness against attacks.